Two search and rescue teams leave base at the same time looking for a lost boy. The first team on foot, heads north at 2mph; and the other, on horseback, head south at 4mph. How long will it take them to search a combined distance of 21 miles between them?

Answer by TimothyLamb(4379)

s = d / t
t = d / s
---
t = 21/(2 + 4)
t = 21/6
---
t = 3.5 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php